# Simpler Diffusion (SiD2): 1.5 FID on ImageNet512 with pixel-space diffusion

### v1

@@ 5 Discussion
In this paper we show how one can train end-to-end diffusion and achieve competitive performance with latent models. In addition, our method outperforms existing end-to-end models by a large margin. Nevertheless, at this moment latent diffusion still has slightly better scaling properties and it remains to be seen whether this gap can be closed any further. Importantly, for many applications it may not be either practical or desired to train a separate autoencoder. Therefore, we believe this paper is an important contribution that allows end-to-end modelling with a single unified approach.
Appendix A Experimental Details
A.1 Power loss
@@
A.2 Estimating Flops
Although FLOPs can be estimated directly using accelerator tooling, it turns out that the most significant operations outweigh all other minor operations. These most significant operations are matrix multiplications (including convolutional layers) and the self-attention dot product. The compute footprint of these operations for our architecture can be calculated using the following equations. Consistent with (Karras et al., 2023), we assume a training step has the computational cost of a forward pass times 3.
@@ A.2 Estimating Flops
return (2 * 3**2) * num_channels ** 2 * blocks * size ** 2 / 1000**3
A.3 A note on autoguidance
Besides applying guidance on intervals (Kynkäänniemi et al., 2024), there is another technique named autoguidance (Karras et al., 2024) which can even produce better FIDs with EDM-XXL (1.2 from 1.4 with guidance intervals). Autoguidance does not use an unconditional model as negative signal. Instead, the negative signal is provided by a worse model, either earlier in training or a small version. Autoguidance increases the hyperparameter space of guidance even further, and we found that this made autoguidance more difficult to tune. For that reason we opted to compare all methods in literature on guidance intervals if available, and also provided the performance of our model with constant guidance to compare with older methods.
A.4 Experimental settings
@@
block_type = [’ResBlock’, ’ResBlock’, ’TransformerBlock’, ’TransformerBlock’],
@@
block_type = [’ResBlock’, ’ResBlock’, ’TransformerBlock’, ’TransformerBlock’]
@@
# for 128^2 resolutions the top-level layer (128 channels) is entirely removed.
To our surprise, a grid search over guidance strength and guidance interval determined that the same settings were optimal for the small and flop heavy variant on ImageNet512. We then shifted the lowerbound of the guidance interval based on resolution shift. The important sampling settings are:
@@
diffusion_schedule = ’cosine_interpolated_low_32_high_512’
@@
max_train_steps = 800_000 # for flop heavy variants
Appendix B Additional Results
5050505010010010010020020020020050050050050010001000100010002000200020002000111122223333555510101010ADMADM-UDiT-XL/2 (interval)U-ViT, LVDM++EDM2-S (interval)EDM2-XXL (interval)SmallSmallFlop HeavyFlop HeavyModel complexity (gigaflops per evaluation), ImageNet-512Previous (latent)PreviousOurs (no interval)Ours
Figure 11: Model forward pass complexity.
@@ Appendix B Additional Results
B.1 Timeshifted Sigmoid Losses
The bias term in the sigmoid loss function could be used to emphasize on the low-level information (bits) of the generated images. One could argue that during the course of training, it is important to first focus on the structure of the image and only later penalise mismatches in the high-frequency details. Therefore, we ran a set of experiments with a time shifted bias, where the bias is interpolated from a starting value to the (fixed) end value (-3) in 100k or 200k training steps. In Figure 13, we report the FID over the course of training. We observe that the time shifted bias could bring a small improvement, albeit resulting in two additional hyper parameters (starting value and number of steps).
From the results we observe that shifting the loss during training generally leads to an improvement. Especially when the bias is not too low: shifting from bias -8/-6 to -3 is a good idea, as long as the warmup is relatively quick (100k iterations). Surprisingly, for warmups to 200k iterations the time shifted models have worse performance in early training of stages.
The time shifted bias is increased with a linear warmup, where we start with bias $b_{\textrm{start}}$ and interpolate over $t_{b}$ steps towards $b_{\textrm{end}}$, yielding: $b_{t}=b_{\textrm{start}}+(b_{\textrm{end}}-b_{\textrm{start}})\cdot\min(t/t_{\textrm{b}},1)$. For these experiment, we fix $b_{\textrm{end}}=-3.0$ and use $t_{\textrm{b}}=200k$ steps, and only vary the starting bias $b_{\textrm{s}}$.
@@ B.1 Timeshifted Sigmoid Losses
B.2 Asymmetric U-ViTs
The results are in Figure 14. Here we show the difference when a different number of encoding and decoding blocks are used. We see that increasing depth helps, and decoder could be a little heavier than the encoder, although symmetric scaling also already works reasonably well.
@@ B.2 Asymmetric U-ViTs
Figure 14: Encoder Decoder Experiments

### v2

@@ 5 Discussion
In this paper we present Simpler Diffusion (SiD2) and show how one can train end-to-end diffusion and achieve competitive performance with latent models: (1) use the sigmoid loss, (2) with a simpler U-ViT architecture, and (3) favour a flop heavy model design. As a result, our method outperforms existing end-to-end models by a large margin. Nevertheless, at this moment latent diffusion methods still seem to have slightly better scaling properties and it remains to be seen whether this gap can be closed any further. Importantly, for many applications training a separate autoencoder may not be either practical or desired. Therefore, we believe this paper is an important contribution that allows end-to-end modelling directly in the pixel-space.
Supplementary Material
Appendix A Connection between sigmoid loss and low-bit training
Weighted diffusion losses can be closely related to low-bit training, as shown in [25]. In earlier normalizing flow work, low-bit training has been shown to achieve higher perceptual quality [24]. The reason that low-bit training improves sample fidelity is quite intuitive: by throwing away less significant (and thus important) bits from the training data, the model spends more capacity on modeling more significant bits. Here we will show how sigmoid loss is related to low-bit training, and therefore provide a theoretical understanding on how sigmoid loss balances the bits from the data of different importance.
Training diffusion models on different bit precisions and their loss contributions may give results that are entangled with optimization hyperparameters. We instead consider a simplified data distribution setting: assume the data is univariate, with a uniform distribution over the $2^{n}$ possible values, where n is the bit precision. We further assume the model to be optimal for all noise levels, which is a mixture-of-Gaussians whose means are the $2^{n}$ possible values and variance is determined by the log signal-to-noise ratio $\lambda$. A diffusion loss over the entire dataset can be expressed as:
$L$ $=\mathbb{E}_{q({\bm{x}})}\mathbb{E}_{\lambda\sim q(\lambda)}w(\lambda)\|{\bm{x}}-\hat{{\bm{x}}}_{\lambda}\|^{2},$ (6)
$=\mathbb{E}_{q({\bm{x}})}\mathbb{E}_{\lambda\sim q(\lambda)}\tilde{w}(\lambda)\|{\bm{{\epsilon}}}-\hat{{\bm{{\epsilon}}}}_{\lambda}\|^{2},$ (7)
$=\mathbb{E}_{\lambda\sim q(\lambda)}\tilde{w}(\lambda)\mathbb{E}_{q({\bm{x}})}\|{\bm{{\epsilon}}}-\hat{{\bm{{\epsilon}}}}_{\lambda}\|^{2},$ (8)
where $q(\lambda)$ is determined by the mapping $\lambda_{t}$ and $t\sim\mathcal{U}(0,1)$, and $q({\bm{x}})$ is the data distribution. Given the simplified data distribution and the optimal model, we can compute $\mathbb{E}_{q({\bm{x}})}\|{\bm{{\epsilon}}}-\hat{{\bm{{\epsilon}}}}_{\lambda}\|^{2}$ analytically. Figure 11 shows its value over $\lambda$ for multiple data distributions of different bit precision $n$.
Figure 11: Expected loss over log signal-to-noise ratio $\lambda$ for multiple data distributions of different bit precision.
We can also plot the weighted loss with weighting function $\tilde{w}(\lambda)$ at different noise levels $\tilde{w}(\lambda)\mathbb{E}_{q({\bm{x}})}\|{\bm{{\epsilon}}}-\hat{{\bm{{\epsilon}}}}_{\lambda}\|^{2}$. Given that the optimal sigmoid weighting at resolution $128$ is $\tilde{w}(\lambda)=\sigma(1-\lambda)$, we follow [14] to shift the weighting to $\tilde{w}(\lambda)=\sigma(10.7-\lambda)$ for the univariate example, where the additional bias term comes from $2\log(128)$. The weighted loss is shown in Figure 12. Interestingly, if we follow the low-bit training but instead of assuming the data is a uniform distribution at a certain precision, we assume it is a mixture of multiple precisions, and then the loss ends up being very similar to the one with the sigmoid weighting (Figure 12). The mixture we visualize here is $[8,7,6,5]$-bit with a mixture proportion of $[1,4,4,6]$. Informally speaking, the loss of the sigmoid weighting is very similar training on a lower-bit distribution of the data, where the data is only modelled at full precision occasionally.
Figure 12: Weighted loss with a sigmoid weighting, compared with a low-bit training of a mixture of precisions.
Appendix B Experimental Details
B.1 Sigmoid Loss
The sigmoid loss can be easily implemented in the following way:
def sigmoid_loss(x, model_x, logsnr_fn, t, bias):
logsnr = logsnr_fn(t)
dlogsnr_dt = jax.jvp(
logsnr_fn, t (ones_like(t),))[1]
weight = -0.5 * dlogsnr_dt * exp(
bias) * sigmoid(logsnr - bias)
return weight * mean_except_batch((x-model_x)**2)
For completeness, recall that a cosine interpolated logsnr schedule function from [14] can be defined as:
def cosine_interpolated(
t, lognsr_min=-10, logsnr_max=10,
image_res=512, noise_res_low=32,
noise_res_high=512)
log_change_high = log(image_res) - log(noise_res_high)
log_change_low = log(image_res) - log(noise_res_low)
b = arctan(exp(-0.5 * logsnr_max))
a = arctan(exp(-0.5 * logsnr_min)) - b
logsnr_cosine = -2. * (jnp.log(jnp.tan(a * t + b))
logsnr_high = logsnr_cosine + log_change_high
logsnr_low = logsnr_cosine + log_change_low
return (1 - t) * logsnr_high + t * logsnr_low
Observe here that the min and max values are shifted along as well and are not the limits of the resulting function.
B.2 Power loss
@@
B.3 Estimating Flops
Although FLOPs can be estimated directly using accelerator tooling, it turns out that the most significant operations outweigh all other minor operations (within 1%). These most significant operations are matrix multiplications (including convolutional layers) and the self-attention dot product. The compute footprint of these operations for our architecture can be calculated using the following equations. Consistent with [20], we assume a training step has the computational cost of a forward pass times 3, and that multiply-adds are counted as one flop.
@@ B.3 Estimating Flops
flops = 2 * 3**2 * blocks # 2 layers with 3x3 conv
flops *= num_channels ** 2 # channels
flops *= size ** 2 # spatial resolution
return flops / 1000**3
B.4 Guidance Intervals
We use guidance intervals [27], in which classifier-free guidance is only applied for noise levels within the interval and disabled elsewhere. Based on a grid search we found that the combination of guidance 1.0 on logsnr (-3, +5) obtained the best FID quality on ImageNet512. Without further tuning we found that the minimum logsnr could simply be shifted by approximately 1.5 for a 2x resolution change, similar to how loss biases are shifted. The logsnr max value is typically less sensitive [27], and was kept constant.
B.5 A note on auto-guidance
Besides applying guidance on intervals [27], there is another technique named autoguidance [21] which can even produce better FIDs with EDM-XXL (1.2 from 1.4 with guidance intervals). Autoguidance does not use an unconditional model as negative signal. Instead, the negative signal is provided by a worse model, either earlier in training or a small version. Autoguidance increases the hyperparameter space of guidance even further, and we found that this made autoguidance more difficult to tune. For that reason we opted to compare all methods in literature on guidance intervals if available, and also provided the performance of our model with constant guidance to compare with older methods.
B.6 Experimental settings
@@
block_type = [’ResBlock’, ’ResBlock’,
’Transformer’, ’Transformer’],
@@
block_type = [’ResBlock’, ’ResBlock’,
’Transformer’, ’Transformer’]
@@
# for 128^2 resolutions the top-level
# layer (128 channels) is entirely removed.
To our surprise, a grid search over guidance strength and guidance interval determined that the same settings were optimal for the small and flop heavy variant on ImageNet512. We then shifted the lower bound of the guidance interval based on resolution shift. Resulting in:
@@
diffusion_schedule =
’cosine_interpolated_low_32_high_512’
@@
max_train_steps = 800_000 # for flop heavy
B.7 Kinetics600
The training settings are:
batch_size=512
optimizer=’adam’,
adam_beta1=0.9
adam_beta2=0.98
adam_eps=1.e-10
learning_rate=1e-4
learning_rate_warmup_steps=100
diffusion_schedule =
’cosine_interpolated_low_32_high_128’
weight_decay=0.0
ema_decay=0.9999
max_train_steps = 500_000
And the model settings are:
channels = [128, 256, 512, 1024]
num_updown_blocks = [3, 3, 4],
num_mid_blocks = 8,
block_dropout = [0., 0., 0.1, 0.1],
block_type = [’ResBlockConv3D’, ’ResBlockConv3D’,
’LocalTransformer_19_19_19’,
’Transformer’],
mean_type = v
loss_type = sigmoid:0
guidance = 0. # unconditional
num_steps = 128
sampler = ’aDDIM’
sampler_noise = ’data_0.9’
clip_x = ’static’
Using the same gflop calculations as before adapted to 3D Convs, our 3570 GFLOPs per forward pass, and about 2.5 zettaflops for training. For reference, the base model of W.A.L.T. uses approximately 0.5 zettaflops (not counting the autoencoder training). This shows that pixel-based models can outperform latent approaches and scale much better using our tuning than before. On the other hand, they still require more compute to achieve this performance that latent approaches. During sampling we use aDDIM [10] because we found that it allowed fewer sampling steps ($128$) with acceptable performance without much noise schedule tuning, making evaluation during training a bit faster.
@@ B.7 Kinetics600
Figure 15: Encoder-Decoder Experiments
5050505010010010010020020020020050050050050010001000100010002000200020002000111122223333555510101010ADMADM-UDiT-XL/2 (interval)U-ViT, LVDM++EDM2-S (interval)EDM2-XXL (interval)SmallSmallFlop HeavyFlop HeavyPrevious (latent)Previous (pixel)Ours (no interval)Ours
Figure 16: Model forward pass complexity (gigaflops) on ImageNet512, figure adapted from [20]. Note: both axis are in log-scale. See also Table 5 for more details.
Table 5: Comparison on sampling cost in FLOPs and generation methods.
Method NFE Inference cost (GFLOPs) Sampling cost (TFLOPs) Sampling time (s) FID
StyleGAN-XL [38] 1 - - 0.10 2.40
EDM2-S [20] 63 102 42.5 - 1.68
EDM2-XXL [20] 63 552 61.6 - 1.40
DiT-XL/2 [31] 250 525 262.5 - 2.40
Pagoda [22] 1 - - 0.05 1.80
SiD2 16-step distilled (ours) 16 653 10.4 0.29 1.50
Appendix C Additional Results
C.1 No Guidance
C.2 Timeshifted Sigmoid Losses
The bias term in the sigmoid loss function could be used to emphasize on the low-level information (bits) of the generated images. One could argue that during the course of training, it is important to first focus on the structure of the image and only later penalise mismatches in the high-frequency details. Therefore, we ran a set of experiments with a time shifted bias, where the bias is increased with a linear warmup, we start with bias $b_{\textrm{start}}$ and interpolate over $t_{b}$ steps towards $b_{\textrm{end}}$, yielding: $b_{t}=b_{\textrm{start}}+(b_{\textrm{end}}-b_{\textrm{start}})\cdot\min(t/t_{\textrm{b}},1)$. So the the bias is interpolated from a starting value to the (fixed) end value (-3) in 100k or 200k training steps.
In Figure 14, we report the FID over the course of training. From the results we observe that shifting the loss during training generally leads to a small improvement, albeit resulting in two additional hyper parameters (starting value and number of steps). Especially when the bias is not too low: shifting from bias -8/-6 to -3 is a good idea, as long as the warmup is relatively quick (100k iterations). Surprisingly, for warmups to 200k iterations the time shifted models have worse performance in early training of stages.
C.3 Asymmetric U-ViTs
In our U-ViT design only a single skip connection per encoding/decoding level is used, allowing for different number of blocks in each level. Here we experiment using different numbers of encoding blocks versus decoding blocks at the same level. The results are in Figure 15. We see that increasing depth helps, and decoder could be a little heavier than the encoder, although symmetric scaling (E3-D3, our baseline) already acts as a strong baseline.
C.4 Forward pass complexity
In addition to the training cost in the main text, Figure 16 shows the forward pass complexity. Here we see that again our method outperforms all existing pixel approaches by a large margin, and performs similar albeit somewhat worse than latent approaches.
C.5 Additional samples
We provide more illustrative examples for ImageNet512 generation (Figure 17) and Text-to-Image (Figure 18 and Figure 19).
Figure 17: Some random (not cherry-picked) class-conditional samples of $512\times 512$ that have been generated by SiD2 flop heavy. Guidance 2.0 on interval logsnr $\in$ (-8, 5). Every row uses the same class-conditioning, every column uses the same rng.
Refer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to caption
Figure 18: Generations from the text-to-image model. Used prompts (the Cartesian product of the sets): "{A strawberry elephant, A corgi wearing a hat, A bird made of seeds, Pixel art: a man walking} in a {lush forest, busy street}."
Refer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to caption
Figure 19: Generations from the distilled text-to-image model. Guidance is set to +3. For each prompt two images selected from four generated images. Used prompts: (1) A couple gets caught in the rain, oil on canvas, (2) A lone traveller walks in a misty forest, (3) A walking figure made out of water, (4) In the swamp, a crocodile stealthily surfaces, revealing only its eyes and the tip of its nose as it moves forward, (5) A fox dressed in suit dancing in park, (6) Pouring chocolate sauce over vanilla ice cream in a cone, studio lighting, (7) An astronaut riding a horse, (8) Aurora Borealis Green Loop Winter Mountain Ridges Northern Lights, (9) Sailboat sailing on a sunny day in a mountain lake, (10) A dog driving a car on a suburban street wearing funny sunglasses.